# Shortcuts and Identifiability in Concept-based Models from a Neuro-Symbolic Lens

### v1

@@ Acknowledgements
The authors are grateful to Antonio Vergari, Emile van Krieken and Luigi Gresele for useful discussions. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA). Neither the European Union nor the granting authority can be held responsible for them. Grant Agreement no. 101120763 - TANGO
@@ Appendix A Implementation Details
Here, we provide additional details about metrics, datasets, and models, for reproducibility. All the experiments were implemented using Python 3.9 and Pytorch 1.13 and run on one A100 GPU. The implementations of DPL, and the code for RSs mitigation strategies were taken verbatim from [Marconato et al., 2023b], while the code for DSL was taken from [Daniele et al., 2023]. DPL and bears were implement on top of DSL codebase.
A.1 Datasets
All data sets were generated using the rsbench library [Bortolotti et al., 2024].
A.1.1 MNIST-Add
MNIST-Add [Manhaeve et al., 2018] consists of pairs of MNIST digits [LeCun, 1998], ranging from 00 to $9$, and the goal is to predict their sum. The prior knowledge used for generating the labels is simply the rule of addition: $\mathsf{K}=(Y=C_{1}+C_{2})$. This task does not admit RSs nor JRSs when digits are processed separately. Below we report four examples:
@@ A.1.1 MNIST-Add
A.2 MNIST-SumParity
@@
[…]
We consider two variants: like MNIST-Add, in MNIST-SumParity proper the training and test splits contain all possible combinations of digits, while in biased MNIST-SumParity the training set contains (even, even), (odd, odd) and (odd, even) pairs, while the test set contains (even, odd) pairs. This fools CBMs into learning an incorrect inference layer implementing subtraction instead of addition-and-parity, as illustrated in Fig. 1. Both variants have the same number of training, validation, and test examples as MNIST-Add.
A.3 Clevr
Clevr [Johnson et al., 2017] consists of of 3D scenes of simple objects rendered with Blender, see Fig. 4. Images are $3\times 128\times 128$ and contain a variable number of objects. We consider only images with $2$ to $4$ objects each, primarily due to scalability issues with DPL. The ground-truth labels were computed by applying the rules (prior knowledge) proposed by [Stammer et al., 2021], reported in Table 6. Objects are entirely determined by four concepts, namely shape, color, material, and size. We list all possible values in Table 5. Overall, Clevr has $6000$ training examples, $1200$ validation examples, and $1800$ test examples.
@@ A.3 Clevr
Preprocessing. All CBMs process objects separately. Following [Shindo et al., 2023], we fine-tune a pretrained YoloV11 model [Jocher et al., 2023] (for $10$ epochs, random seed $13$) on a subset of the training set’s bounding boxes, and use it to predict bounding boxes for all images. For each training and test image, we concatenate the regions in the bounding boxes, obtaining a list $2$ to $4$ images which plays the role of input for our CBMs.
@@
A.4 Metrics
@@
Label quality. We measure the macro average $F_{1}$-score and Accuracy, to account for class imbalance.
Concept quality. We compute concept quality via Accuracy and $F_{1}$-score. However, the order of ground-truth and predicted concepts might might differ, e.g., in Clevr $G_{1}$ might represent whether “color = red”, while $C_{1}$ whether “color = blue”. Therefore, in order to compute these metrics we have to first align them using the Hungarian matching algorithm using concept-wise Pearson correlation on the test data [Kuhn, 1955]. The algorithm computes the correlation matrix $R$ between $\bm{\mathrm{g}}$ and $\bm{\mathrm{c}}$, i.e., $R_{ij}=\mathrm{corr}(G_{i},C_{j})$ where $\mathrm{corr}$ is the correlation coefficient, and then infers a permutation $\pi$ between concept values (e.g., mapping ground-truth color “red” to predicted value “blue”) that maximizes the above objective. The metrics are computed by comparing the reordered ground-truth annotations with the predicted concepts.
@@
For MNIST-Add and MNIST-SumParity, we conducted an exhaustive evaluation since the number of possible combinations to supervise is $100$. The results in Table 1 and Table 2 evaluate the knowledge exhaustively, whereas in Table 4, we separately assess the knowledge of in-distribution and out-of-distribution combinations.
In Clevr, as discussed in Section A.3, there are too many combinations of concept values to possibly evaluate them all. Therefore, both supervision and evaluation are performed by randomly sampling $100$ combinations of ground-truth concepts.
A.5 Architectures
For MNIST-SumParity and MNIST-Add, We employed the architectures from [Bortolotti et al., 2024], while for DSL, DPL, and bears, we followed the architecture described in [Daniele et al., 2023]. For Clevr instead, we adopted the decoder presented in Table 7 and the encoder presented in Table 8. All models process different objects (e.g., digits) separately using the same backbone.
@@ A.5 Architectures
Details of SENN. SENN [Alvarez-Melis and Jaakkola, 2018] are unsupervised concept-based models consisting of a neural network that produces a distribution over concepts, a neural network that assigns a score to those concepts, and a decoder. The final prediction is made by computing the dot product between the predicted concepts and their predicted scores. SENN by default includes a reconstruction penalty, a robustness loss, and a sparsity term.
@@
Details of bears. bears combines DSL [Daniele et al., 2023] and BEARS [Marconato et al., 2024]. BEARS is a NeSy architectures based on deep ensembles [Lakshminarayanan et al., 2017], designed to improve the calibration of NeSy models in the presence of RSs and provide uncertainty estimation for concepts that may be affected by such RSs and it assumes to be given prior knowledge.
@@ A.6 Mitigation Strategies
Entropy maximization. The entropy regularization term follows the implementation in [Marconato et al., 2023b]. The loss is defined as $1-\frac{1}{k}\sum_{i=1}^{k}H_{m_{i}}[p_{\theta}(c_{i})]$ where $p_{\theta}(C)$ represents the marginal distribution over the concepts, and $H_{m_{i}}$ is the normalized Shannon entropy over $m_{i}$ possible values of the distribution.
@@
Reconstruction. The reconstruction penalty is the same as in [Marconato et al., 2023b]. In this case, the concept extractor outputs both concepts $\bm{\mathrm{c}}$ and style variables $\bm{\mathrm{z}}$, i.e., it implements a distribution $p_{\theta}(c,z|x)=p_{\theta}(c|x)\cdot p_{\theta}(z|x)$. The reconstruction penalty is defined as: $-\mathbb{E}_{(c,z)\sim p_{\theta}(c,z|x)}\left[\log p_{\psi}(x|c,z)\right]$ where $p_{\psi}(x|c,z)$ represents the decoder network.
@@
Contrastive loss. We implemented contrastive learning using the InfoNCE loss, comparing the predicted concept distribution of the current batch with that of its respective augmentations. Following the recommendations of Chen et al. [2020a], we applied the following augmentations to each batch: random cropping (20 pixels for MNIST and 125 pixels for Clevr), followed by resizing, color jittering, and Gaussian blur.
A.7 Model Selection
@@
Optimization. For DPL, bears, CBNM and SENN we used the Adam optimizer [Kingma, 2014], while for DSL and DPL we achieved better results using the Madgrad optimizer [Defazio and Jelassi, 2021].
@@
Hyperparameter search. We performed an extensive grid search on the validation set over the following hyperparameters: (i) Learning rate ($\gamma$) in {1e-4, 1e-3, 1e-2}; (ii) Weight decay ($\omega$) in {0, 1e-4, 1e-3, 1e-2, 1e-1}; (iii) Reconstruction, contrastive loss, entropy loss, concept supervision loss and knowledge supervision loss weights ($w_{r}$, $w_{c}$, $w_{h}$, $w_{csup}$ and $w_{k}$, respectively) in {$0.1$, $1$, $2$, $5$, $8$, $10$}; (iv) Batch size ($\nu$) in {$32$, $64$, $128$, $256$, $512$}. DSL and DPL have additional hyperparameters for the truth table: $\varepsilon_{sym}$ and $\varepsilon_{rul}$ for DSL, and $\lambda_{r}$, the truth table learning rate, for DPL.
@@
All experiments were run for approximately $50$ epochs for MNIST variants and $100$ epochs for Clevr using early stopping based on validation set $\mathrm{F_{1}}(Y)$ performance. The exponential decay rate $\beta$ was set to $0.99$ for all experiments, as we empirically found it to provide the best performance across tasks. In all experiments, we selected $\gamma=1e-3$.
@@
For the second research question, we performed a grid search for each mitigation strategy individually. Additionally, since our analysis focuses on optimal models, we trained multiple models and retained only those achieving optimal $\mathrm{F_{1}}(Y)$ performance on the validation set. For MNIST-SumParity, we set $\nu=128$, $\omega=0.001$, and $w_{csup}=1$ across all experiments. For specific mitigation strategies, we used $w_{h}=1$ for entropy regularization, $w_{r}=0.01$ for reconstruction, $w_{c}=0.1$ for contrastive learning, and $w_{k}=1$ for knowledge supervision. For Clevr, we applied the same settings: $\nu=128$, $\omega=0.001$, and $w_{csup}=1$ for all experiments. The specific mitigation strategy weights were set as follows: $w_{h}=1$ for entropy regularization, $w_{r}=1$ for reconstruction, $w_{c}=0.1$ for contrastive learning, and $w_{k}=1$ for knowledge supervision. When concept supervision is applied, we follow a sequential training approach as in [Koh et al., 2020]. During the initial epochs, the model learns only the concepts, and later, it jointly optimizes both the concepts and the label, as they are challenging to learn simultaneously.
Appendix B Additional results
B.1 Full Results for the Case Studies
We report additional results for Q1 including additional models and metrics – namely label accuracy $\mathrm{Acc}_{Y}$, concept accuracy $\mathrm{Acc}_{C}$, and inference layer accuracy $\mathrm{Acc}(\bm{\mathrm{\beta}})$, and negative log-likelihood on the test set. The two additional models are: “CBNM noent”, like CBNM except it does not include any entropy term; and CBNM 20 which is a regular CBNM for which $20\%$ of the concept values (e.g., two digits in MNIST-based tasks) receive full concept supervision.
Table 9: Complete results for MNIST-Add.
Method $\mathrm{Acc}_{Y}$ $\mathrm{F_{1}}(Y)$ $\mathrm{Acc}_{C}$ $\mathrm{F_{1}}(C)$ $\mathsf{Cls}(C)$ ($\downarrow$) $\mathrm{F_{1}}(\bm{\mathrm{\beta}})$ $\mathrm{Acc}(\bm{\mathrm{\beta}})$ $\mathsf{NLL}$
DPL $0.98\pm 0.01$ $0.98\pm 0.01$ $0.99\pm 0.01$ $0.99\pm 0.01$ $0.01\pm 0.01$- -- $0.17\pm 0.03$
CBNM $0.98\pm 0.01$ $0.98\pm 0.01$ $0.99\pm 0.01$ $0.99\pm 0.01$ $0.01\pm 0.01$ $1.00\pm 0.01$ $1.00\pm 0.01$ $0.11\pm 0.01$
CBNM noent $0.98\pm 0.01$ $0.98\pm 0.01$ $0.80\pm 0.10$ $0.74\pm 0.13$ $0.08\pm 0.04$ $0.37\pm 0.09$ $0.42\pm 0.06$ $0.26\pm 0.01$
CBNM 20 $0.98\pm 0.01$ $0.98\pm 0.01$ $0.99\pm 0.01$ $0.99\pm 0.01$ $0.01\pm 0.01$ $0.67\pm 0.02$ $0.59\pm 0.02$ $0.70\pm 0.02$
CBNM 20 noent $0.92\pm 0.02$ $0.94\pm 0.01$ $0.60\pm 0.09$ $0.48\pm 0.10$ $0.22\pm 0.12$ $0.02\pm 0.01$ $0.12\pm 0.02$ $1.09\pm 0.01$
DSL $0.96\pm 0.02$ $0.96\pm 0.02$ $0.98\pm 0.01$ $0.98\pm 0.01$ $0.01\pm 0.01$ $1.00\pm 0.01$ $1.00\pm 0.01$ $0.38\pm 0.14$
DPL $0.90\pm 0.09$ $0.91\pm 0.09$ $0.93\pm 0.07$ $0.92\pm 0.08$ $0.01\pm 0.01$ $0.90\pm 0.08$ $0.90\pm 0.09$ $0.02\pm 0.01$
bears $0.81\pm 0.12$ $0.76\pm 0.15$ $0.87\pm 0.12$ $0.87\pm 0.13$ $0.01\pm 0.01$ $0.67\pm 0.14$ $0.65\pm 0.10$ $0.34\pm 0.22$
SENN $0.97\pm 0.01$ $0.97\pm 0.01$ $0.84\pm 0.05$ $0.80\pm 0.07$ $0.01\pm 0.01$ $0.75\pm 0.08$ $0.75\pm 0.08$ $0.01\pm 0.01$
@@ B.1 Full Results for the Case Studies
Method $\mathrm{Acc}_{Y}$ $\mathrm{F_{1}}(Y)$ $\mathrm{Acc}_{C}$ $\mathrm{F_{1}}(C)$ $\mathsf{Cls}(C)$ ($\downarrow$) $\mathrm{F_{1}}(\bm{\mathrm{\beta}})$ $\mathrm{Acc}(\bm{\mathrm{\beta}})$ $\mathsf{NLL}$
DPL $0.99\pm 0.01$ $0.99\pm 0.01$ $0.47\pm 0.04$ $0.43\pm 0.08$ $0.36\pm 0.15$- -- $0.33\pm 0.01$
CBNM $0.90\pm 0.18$ $0.90\pm 0.18$ $0.21\pm 0.06$ $0.09\pm 0.03$ $0.66\pm 0.09$ $0.54\pm 0.04$ $0.21\pm 0.06$ $0.52\pm 1.04$
CBNM noent $0.98\pm 0.01$ $0.98\pm 0.01$ $0.22\pm 0.01$ $0.08\pm 0.02$ $0.72\pm 0.13$ $0.53\pm 0.09$ $0.53\pm 0.09$ $0.07\pm 0.02$
DSL $0.94\pm 0.03$ $0.94\pm 0.03$ $0.20\pm 0.01$ $0.07\pm 0.01$ $0.80\pm 0.01$ $0.52\pm 0.01$ $0.51\pm 0.02$ $0.29\pm 0.16$
DPL $0.99\pm 0.01$ $0.99\pm 0.01$ $0.22\pm 0.01$ $0.07\pm 0.01$ $0.80\pm 0.01$ $0.50\pm 0.05$ $0.50\pm 0.06$ $0.06\pm 0.01$
bears $0.99\pm 0.01$ $0.99\pm 0.01$ $0.28\pm 0.05$ $0.30\pm 0.03$ $0.22\pm 0.04$ $0.36\pm 0.09$ $0.37\pm 0.09$ $0.04\pm 0.01$
SENN $0.99\pm 0.01$ $0.99\pm 0.01$ $0.53\pm 0.05$ $0.49\pm 0.05$ $0.01\pm 0.01$ $0.53\pm 0.06$ $0.53\pm 0.06$ $0.01\pm 0.01$

### v2

@@ Acknowledgements
The authors are grateful to Antonio Vergari, Emile van Krieken, Marco Pacini, and Luigi Gresele for useful discussions. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA). Neither the European Union nor the granting authority can be held responsible for them. Grant Agreement no. 101120763 - TANGO. PM was supported by the MSCA project “Probabilistic Formal Verification for Provably Trustworthy AI - PFV-4-PTAI” under GA no. 101110960.
@@ Appendix A Implementation Details
Here, we provide additional details about metrics, datasets, and models, for reproducibility. All the experiments were implemented using Python 3.9 and Pytorch 1.13 and run on one A100 GPU. The implementations of DPL, and the code for RSs mitigation strategies were taken verbatim from [19], while the code for DSL was taken from [50]. DPL and bears were implement on top of DSL codebase.
A.1 Notation
We summarize the notation used in this paper.
Table 6: Glossary of common symbols.
Symbol Meaning
$x$, $y$, $z$ Scalar constants
$X$, $Y$, $Z$ Scalar variables
$\bm{\mathrm{x}}$, $\bm{\mathrm{X}}$ Vector constant and vector variable
$\mathcal{X}$ Space of input
$\mathcal{Y}$, $\mathcal{G}$, $\mathcal{C}$ Sets of labels, ground-truth, and learned concepts
$\Delta_{\mathcal{Y}},\Delta_{\mathcal{G}},\Delta_{\mathcal{C}}$ Simplices of probability distributions on sets elements
$\bm{\mathrm{f}}:\mathcal{X}\to\Delta_{\mathcal{C}}$ Concept extractor
$\bm{\mathrm{\omega}}:\Delta_{\mathcal{C}}\to\Delta_{\mathcal{Y}}$ Inference layer
$\mathcal{F},\Omega$ Space of concept extractors and inference layers
$\bm{\mathrm{\alpha}}:\mathcal{G}\to\Delta_{\mathcal{C}}$ Map from ground-truth concepts to distribution of model concepts
$\bm{\mathrm{\beta}}:\mathcal{C}\to\Delta_{\mathcal{Y}}$ Map of the inference layer on concept vectors
$\mathcal{A},\mathcal{B}$ Spaces of $\bm{\mathrm{\alpha}}$’s and $\bm{\mathrm{\beta}}$’s
$p^{*}$ distribution over inputs, label, and ground-truth concepts
$\mathsf{K}$ Prior knowledge
$\models$ Logical entailment
A.2 Datasets
All data sets were generated using the rsbench library [22].
A.2.1 MNIST-Add
MNIST-Add [41] consists of pairs of MNIST digits [45], ranging from 0 to $9$, and the goal is to predict their sum. The prior knowledge used for generating the labels is simply the rule of addition: $\mathsf{K}=(Y=C_{1}+C_{2})$. This task does not admit RSs nor JRSs when digits are processed separately. The training and test examples take the following form:
@@ A.2.1 MNIST-Add
A.3 MNIST-SumParity
Similarly, in MNIST-SumParity the goal is to predict the parity of the sum, i.e., the prior knowledge is the rule: $\mathsf{K}=(Y=(C_{1}+C_{2})\ \mathrm{mod}\ 2)$. This task is more challenging for our purposes, as it is affected by both RSs and JRSs, cf. Fig.˜1. Below we report four examples:
@@ A.3 MNIST-SumParity
We consider two variants: like MNIST-Add, in MNIST-SumParity proper the training and test splits contain all possible combinations of digits, while in biased MNIST-SumParity the training set contains (even, even), (odd, odd) and (odd, even) pairs, while the test set contains (even, odd) pairs only. This fools CBMs into learning an incorrect inference layer implementing subtraction instead of addition-and-parity, as illustrated in Fig.˜1. Both variants have the same number of training, validation, and test examples as MNIST-Add.
A.4 Clevr
Clevr [83] consists of of 3D scenes of simple objects rendered with Blender, see Fig.˜4. Images are $3\times 128\times 128$ and contain a variable number of objects. We consider only images with $2$ to $4$ objects each, primarily due to scalability issues with DPL. The ground-truth labels were computed by applying the rules (prior knowledge) proposed by [14], reported in Table˜8. Objects are entirely determined by four concepts, namely shape, color, material, and size. We list all possible values in Table˜7. Overall, Clevr has $6000$ training examples, $1200$ validation examples, and $1800$ test examples.
@@ A.4 Clevr
Preprocessing. All CBMs process objects separately, as follows. Following [101], we fine-tune a pretrained YoloV11 model [74] (for $10$ epochs, random seed $13$) on a subset of the training set’s bounding boxes, and use it to predict bounding boxes for all images. For each training and test image, we concatenate the regions in the bounding boxes, obtaining a list $2$ to $4$ images which plays the role of input for our CBMs.
@@
A.5 BDD-OIA
BDD-OIA is an autonomous driving dataset, composed of images extracted from driving scene videos [84]. Each image is annotated with four binary labels: move_forward, stop, turn_left, and turn_right. Each image comes with 21 binary concepts that constitute the explanations for the corresponding actions. The training set contains almost 1616k fully labeled frames, while the validation and test sets include almost 22k and 4.54.5k annotated samples, respectively. We adopt the same prior knowledge used in [22], which we briefly summarize here for the sake of completeness.
For the move_forward and stop actions, the rules are as follows:
[…]
The rules for the turn_left and the turn_right action, instead, are:
[…]
An overview of the classes of BDD-OIA can be found in [84].
Preprocessing. BDD-OIA images are preprocessed following the approach described in [27]. We first used a Faster R-CNN [102] model pre-trained on MS-COCO and fine-tuned on BDD-100k, then we extract $2048$-dimensional features using a pre-trained convolutional layer from [27].
A.6 Metrics
@@
Label quality. We measure the macro average $F_{1}$-score and Accuracy, to account for imbalance.
Concept quality. We compute concept quality via Accuracy and $F_{1}$-score. However, the order of ground-truth and predicted concepts might might differ, e.g., in Clevr $G_{1}$ might represent whether “color = red”, while $C_{1}$ whether “color = blue”. Therefore, in order to compute these metrics we have to first align them using the Hungarian matching algorithm using concept-wise Pearson correlation on the test data [64]. The algorithm computes the correlation matrix $R$ between $\bm{\mathrm{g}}$ and $\bm{\mathrm{c}}$, i.e., $R_{ij}=\mathrm{corr}(G_{i},C_{j})$ where $\mathrm{corr}$ is the correlation coefficient, and then infers a permutation $\pi$ between concept values (e.g., mapping ground-truth color “red” to predicted value “blue”) that maximizes the above objective. The metrics are computed by comparing the reordered ground-truth annotations with the predicted concepts.
Concept Collapse.
To compute the Concept Collapse metric $\mathsf{Cls}(C)$ [22], we follow the procedure described in Bortolotti et al. [22], which we briefly describe here. Specifically, we extract a confusion matrix in the multilabel setting by encoding each binary concept vector (e.g., $(0,1,1)\mapsto 3$) into a categorical label. Let $\mathcal{C}^{*}\subseteq\{0,1\}^{k}$ be the set of ground-truth concept vectors in the test set, and $\mathcal{C}$ be the set of predicted concept vectors. After converting both sets to categorical values, denoted $\mathcal{F}(\mathcal{C}^{*})$ and $\mathcal{F}(\mathcal{C})$ respectively, we construct the confusion matrix $C\in[0,1]^{m\times m}$, where $m=|\mathcal{F}(\mathcal{C})\cup\mathcal{F}(\mathcal{C}^{*})|$. From this matrix, we compute the number of predicted categories $p$ as:
$p=\sum_{j=1}^{m}\mathbbm{1}\!\{\exists i\ .\ C_{ij}>0\}=|\mathcal{F}(\mathcal{C})|$ (12)
Then, the $\mathsf{Cls}(C)$ metric is given by:
$\mathsf{Cls}(C)=1-\frac{p}{m}$ (13)
Where $m=2^{k}$.
@@
For MNIST-Add and MNIST-SumParity, we conducted an exhaustive evaluation since the number of possible combinations to supervise is $100$. The results in Table˜5 and Table˜5 evaluate the knowledge exhaustively, whereas in Table˜5, we separately assess the knowledge of in-distribution and out-of-distribution combinations.
In Clevr, as discussed in Section˜A.4, there are too many combinations of concept values to possibly evaluate them all. Therefore, both supervision and evaluation are performed by randomly sampling $100$ combinations of ground-truth concepts. We follow the same procedure for BDD-OIA, except that we sample more combinations ($2000$) due to the larger space of possible concept configurations.
A.7 Architectures
For MNIST-SumParity and MNIST-Add, We employed the architectures from [22], while for DSL, DPL, and bears, we followed the architecture described in [50]. For Clevr instead, we adopted the decoder presented in Table˜9 and the encoder presented in Table˜10. All models process different objects (e.g., digits) separately using the same backbone. For BDD-OIA, we adopt the architectures defined in [22].
@@ A.7 Architectures
Details of SENN. SENN [1] are unsupervised concept-based models consisting of a neural network that produces a distribution over concepts, a neural network that assigns a score to those concepts, and a decoder. The final prediction is made by computing the dot product between the predicted concepts and their predicted scores. SENN by default includes a reconstruction penalty, a robustness loss, and a sparsity term.
@@
Details of bears. bears combines DSL [50] and BEARS [63]. BEARS is a NeSy architectures based on deep ensembles [103], designed to improve the calibration of NeSy models in the presence of RSs and provide uncertainty estimation for concepts that may be affected by such RSs and it assumes to be given prior knowledge.
@@ A.8 Mitigation Strategies
Entropy maximization. The entropy regularization term follows the implementation in [19]. The loss is defined as $1-\frac{1}{k}\sum_{i=1}^{k}H_{m_{i}}[p_{\theta}(c_{i})]$ where $p_{\theta}(C)$ represents the marginal distribution over the concepts, and $H_{m_{i}}$ is the normalized Shannon entropy over $m_{i}$ possible values of the distribution.
@@
Reconstruction. The reconstruction penalty is the same as in [19]. In this case, the concept extractor outputs both concepts $\bm{\mathrm{c}}$ and style variables $\bm{\mathrm{z}}$, i.e., it implements a distribution $p_{\theta}(c,z\mid x)=p_{\theta}(c\mid x)\cdot p_{\theta}(z\mid x)$. The reconstruction penalty is defined as: $\mathbb{E}_{(c,z)\sim p_{\theta}(c,z\mid x)}\left[\log p_{\psi}(x\mid c,z)\right]$ where $p_{\psi}(x\mid c,z)$ represents the decoder network.
@@
Contrastive loss. We implemented contrastive learning using the InfoNCE loss, comparing the predicted concept distribution of the current batch with that of its respective augmentations. Following the recommendations of Chen et al. [80], we applied the following augmentations to each batch: random cropping (20 pixels for MNIST and 125 pixels for Clevr), followed by resizing, color jittering, and Gaussian blur.
A.9 Model Selection
@@
Optimization. For DPL, bears, CBNM and SENN we used the Adam optimizer [104], while for DSL and DPL we achieved better results using the Madgrad optimizer [105].
@@
Hyperparameter search. We performed an extensive grid search on the validation set over the following hyperparameters: (i) Learning rate ($\gamma$) in {1e-4, 1e-3, 1e-2}; (ii) Weight decay ($\omega$) in {0, 1e-4, 1e-3, 1e-2, 1e-1}; (iii) Reconstruction, contrastive loss, entropy loss, concept supervision loss and knowledge supervision loss weights ($w_{r}$, $w_{c}$, $w_{h}$, $w_{csup}$ and $w_{k}$, respectively) in {$0.1$, 11, 22, 55, 88, 1010}; (iv) Batch size ($\nu$) in {$32$, $64$, $128$, $256$, 512512}. DSL and DPL have additional hyperparameters for the truth table: $\varepsilon_{sym}$ and $\varepsilon_{rul}$ for DSL, and $\lambda_{r}$, the truth table learning rate, for DPL.
@@
All experiments were run for approximately $50$ epochs for MNIST variants, $30$ for BDD-OIA and $100$ epochs for Clevr using early stopping based on validation set $F_{1}(Y)$ performance. The exponential decay rate $\beta$ was set to $0.99$ for all experiments, as we empirically found it to provide the best performance across tasks. In all experiments, we selected $\gamma=1e-3$.
To answer the first research question in Table˜5 on MNIST-Add, for DPL we used: $\nu=32$ and $\omega=0.0001$; for DSL: $\nu=128$, $\omega=0.001$, $\varepsilon_{sym}=0.2807344052335263$, and $\varepsilon_{rul}=0.107711951632426$; for CBNM: $\nu=256$ and $\omega=0.0001$; for DPL: $\nu=32$, $\omega=0.01$, and $\lambda_{r}=0.0001$; for SENN: $\nu=64$, $\omega=0.001$, and $w_{r}=0.1$; while for bears we set the diversification loss term to 0 and the entropy term to $0.1$. On MNIST-SumParity and its biased version, for DPL we used: $\nu=512$ and $\omega=0.0001$; for DSL: $\nu=128$, $\omega=0.0001$, $\varepsilon_{sym}=0.2807344052335263$, and $\varepsilon_{rul}=0.107711951632426$; for CBNM: $\nu=256$ and $\omega=0.0001$; for DPL: $\nu=32$, $\omega=0.01$, and $\lambda_{r}=0.01$; for SENN: $\nu=64$, $\omega=0.001$, and $w_{r}=0.1$; while for bears we set the diversification loss term to $0.1$ and the entropy term to $0.2$. On Clevr, for DPL, CBNM, and DPL, we used $\nu=32$ and $\omega=0.001$; for DPL, $\lambda_{r}=0.001$. On BDD-OIA, for DPL we used $\nu=128$ and $\omega=10^{-4}$, while for CBNM we used $\nu=64$ and $\omega=0$.
For the second research question, we performed a grid search for each mitigation strategy individually. Additionally, since our analysis focuses on optimal models, we trained multiple models and retained only those achieving optimal $F_{1}(Y)$ performance on the validation set. For MNIST-SumParity, we set $\nu=128$, $\omega=0.001$, and $w_{csup}=1$ across all experiments. For specific mitigation strategies, we used $w_{h}=1$ for entropy regularization, $w_{r}=0.01$ for reconstruction, $w_{c}=0.1$ for contrastive learning, and $w_{k}=1$ for knowledge supervision. For Clevr, we applied the same settings: $\nu=128$, $\omega=0.001$, and $w_{csup}=1$ for all experiments. The specific mitigation strategy weights were set as follows: $w_{h}=1$ for entropy regularization, $w_{r}=1$ for reconstruction, $w_{c}=0.1$ for contrastive learning, and $w_{k}=1$ for knowledge supervision. When concept supervision is applied, we follow a sequential training approach as in [3]. During the initial epochs, the model learns only the concepts, and later, it jointly optimizes both the concepts and the label, as they are challenging to learn simultaneously.
Appendix B Additional results
B.1 Additional Results for Q1
Tables˜11, 12 and 13 report additional results for Q1 including two additional models and metrics. The additional models are: “CBNM noent”, which is identical to CBNM except it does not include any entropy term; and CBNM 20 which is a regular CBNM for which $20\%$ of the concept combinations (e.g., pairs of digits in MNIST-based tasks) receive full supervision. The metrics include label accuracy $\mathrm{Acc}_{Y}$, concept accuracy $\mathrm{Acc}_{C}$, inference layer accuracy $\mathrm{Acc}(\bm{\mathrm{\beta}})$, and negative log-likelihood, all evaluated on the test set.
Table 11: Complete results for MNIST-Add.
Method $\mathrm{Acc}_{Y}$ $F_{1}(Y)$ $\mathrm{Acc}_{C}$ $F_{1}(C)$ $\mathsf{Cls}(C)$ ($\downarrow$) $F_{1}(\bm{\mathrm{\beta}})$ $\mathrm{Acc}(\bm{\mathrm{\beta}})$ $\mathsf{NLL}$
DPL $0.98\pm 0.01$ $8\pm 0.01 0.99±0$$9\pm 0.01$ $99\pm 0.01 0.0$$01\pm 0.01 - -$ $0.17\pm 0.03$
@@ B.1 Additional Results for Q1
CBNM noent $0.98\pm 0.01$ $0.98\pm 0.01$ $0.80\pm 0.10$ $74\pm 0.13 0.0$$08\pm 0.04$ $0.37\pm 0.09$ $0.42\pm 0.06$ $0.26\pm 0.01$
[…]
CBNM 20 noent $0.92\pm 0.02$ $0.94\pm 0.01$ $0.60\pm 0.09$ $0.48\pm 0.10$ $22\pm 0.12 0.0$$02\pm 0.01$ $0.12\pm 0.02$ $1.09\pm 0.01$
[…]
DPL $0.90\pm 0.09$ $0.91\pm 0.09$ $0.93\pm 0.07$ $92\pm 0.08 0.0$$01\pm 0.01$ $\pm 0.08 0.90±0.09$$\pm 0.09$ $0.02\pm 0.01$
[…]
SENN $0.97\pm 0.01$ $0.97\pm 0.01$ $0.84\pm 0.05$ $80\pm 0.07 0.0$$01\pm 0.01$ $0.75\pm 0.08$ $75\pm 0.08 0.0$$01\pm 0.01$
@@
Method $\mathrm{Acc}_{Y}$ $F_{1}(Y)$ $\mathrm{Acc}_{C}$ $F_{1}(C)$ $\mathsf{Cls}(C)$ ($\downarrow$) $F_{1}(\bm{\mathrm{\beta}})$ $\mathrm{Acc}(\bm{\mathrm{\beta}})$ $\mathsf{NLL}$
DPL $9\pm 0.01 0.99±0$$9\pm 0.01$ $0.47\pm 0.04$ $0.43\pm 0.08$ $6\pm 0.15 - - 0.33±0$$3\pm 0.01$
[…]
CBNM noent $0.98\pm 0.01$ $0.98\pm 0.01$ $22\pm 0.01 0.0$$08\pm 0.02$ $0.72\pm 0.13$ $0.53\pm 0.09$ $53\pm 0.09 0.0$$07\pm 0.02$
[…]
DPL $9\pm 0.01 0.99±0$$9\pm 0.01$ $22\pm 0.01 0.0$$07\pm 0.01$ $0.80\pm 0.01$ $0.50\pm 0.05$ $50\pm 0.06 0.0$$06\pm 0.01$
[…]
SENN $9\pm 0.01 0.99±0$$9\pm 0.01$ $0.53\pm 0.05$ $49\pm 0.05 0.0$$01\pm 0.01$ $0.53\pm 0.06$ $53\pm 0.06 0.0$$01\pm 0.01$